OpenStudy (anonymous):

WILL GIVE A MEDAAAAL Rewrite the equation below into standard form. 2y=3x+12 A. -3x+2y=12 B. 3x-2y=12 C. 3x-2y=-12 D. x-y=-2

OpenStudy (anonymous):

Standard form is: Ax + By = C where A, B, and C are integers and A is positive.

OpenStudy (anonymous):

So, take the original equation and subtract 2y from each side: 2y - 2y = 3x + 12 - 2y -> 0 = 3x - 2y + 12 Now, subtract 12 from both sides: 0 - 12 = 3x - 2y + 12 - 12 -> -12 = 3x - 2y Now switch sides: 3x - 2y = -12

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @cynthia_franco

OpenStudy (anonymous):

D'awh, no need to thank me, thank you! @tcarroll010

OpenStudy (anonymous):

That's ok, have a great day!